# OpenFly: A Versatile Toolchain and Large-scale Benchmark for Aerial Vision-Language Navigation

### v4

@@ III OpenFly Data Generation Platform
In this section, we first describe several basic simulators and data resources, and then present the developed toolchain. The framework of the whole automatic data generation platform is illustrated in Fig. 2.
III-A Basic Simulators and Data Resources
To collect a wide range of high-quality and realistic simulation data, we source the dataset from multiple rendering engines integrated with various simulators. Fig. 3 showcases several examples obtained from these rendering engines/techniques.
Unreal Engine + AirSim/UnrealCV. UE is a rendering engine capable of providing highly realistic interactive virtual environments. This platform has undergone five iterations, and each version features comprehensive and high-quality digital assets. In UE5, we meticulously select an official sample project named ‘City Sample’, which provides us with a large urban scene covering $25.3km^{2}$ and a smaller one covering $2.7km^{2}$. These scenes include a variety of assets such as buildings, streets, traffic lights, vehicles, and pedestrians. Besides, in UE4, we prepare six more high-quality scenes. Specifically, there are two large scenes showcasing the central urban areas of Shanghai and Guangzhou, covering areas of $30.88km^{2}$ and $58.56km^{2}$, respectively. The remaining four scenes are selected from AerialVLN [29]. They have smaller areas for totally about $26.64km^{2}$. These scenes encompass a wide range of architectural styles, including both Chinese and Western influences, as well as classical and modern designs. Additionally, the UE4 engine allows us to make adjustments in scene time to achieve different appearances of scenes under varying lighting conditions.
Airsim is an open-source simulator, which provides highly realistic simulated environments for UAVs and cars. We integrate the AirSim plugin into UE4 to obtain image data easily from the perspective of a UAV. Since AirSim does not support UE5 and stopped updating in 2022, we use the UnrealCV [53] plugin as an alternative for image acquisition in UE5. To realize a highly efficient data collection in simulated scenes, we modify the UE5 project to a C++ project, integrate the UnrealCV plugin, and package executables for multiple systems like Windows and Linux.
GTA V + Script Hook V. GTA V is an open-world game that is frequently used by computer vision researchers due to its highly realistic and dynamic virtual environment. The game features a meticulously crafted cityscape modeled after Los Angeles, encompassing various buildings and locations such as skyscrapers, gas stations, parks, and plazas, along with dynamic traffic flows and changes in lighting and shadows.
Google Earth + Google Earth Studio. Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery, aerial photographs, and Geographic Information System (GIS) data. From this engine, we select four urban scenes covering a total area of $53.60km^{2}$, i.e., Berkeley, primarily consisting of traditional neighborhoods; Osaka, which features a mix of skyscrapers and historic buildings; and two areas with numerous landmarks: Washington, D.C., and St. Louis.
3D Gaussian Splatting + SIBR viewers. As a highly realistic reconstruction method, hierarchical 3D GS  [21] employs a hierarchical training and display architecture, making it particularly suitable for rendering large-scale areas. Due to these features, we use this method to reconstruct and render multiple real scenes. We utilize the DJI M30T drone as the data collection device, which offers an automated oblique photography mode, enabling us to capture a large area of real-world data with minimal manpower. Practically, we gathered data from five campuses across three universities, which are East China University of Science and Technology, Northwestern Polytechnical University, and Shanghai Jiao Tong University (referred to as ECUST, NWPU, and SJTU). These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns. The detailed information for the five campuses is presented in Table I. More details of the 3D GS data collection can be found in our supplementary material.
TABLE I: Different 3D GS Scenes
@@ III-A Basic Simulators and Data Resources
III-B Toolchain for Automatic Data Colleciton
To achieve automatic data generation, we integrate the above five simulators and design three unified interfaces, i.e., the agent movement interface, the lidar data acquisition interface, and the image acquisition interface, allowing an agent to interact with any scene. Based on these interfaces, we further develop a toolchain, including 3D point cloud acquisition, scene semantic segmentation, automatic trajectory generation, and instruction generation. The framework of the whole data generation platform is illustrated in Fig. 2, with details of these interfaces and tools elaborated below.
@@ III-B Toolchain for Automatic Data Colleciton
Unified Interfaces. 1) Agent Movement Interface: We design a CoorTrans module, which implements a customized pose transformation matrix and scaling function to unify all simulator coordinate systems into a meter-based FLU (Front-Left-Up) convention. This interface enables precise agent positioning among regular scenes, point clouds, and scene segmentations, ensuring consistency and facilitating automatic trajectory generation. 2) Lidar Data Acquisition: For different simulators, point cloud data is acquired through different methods, including lidar sensor collection, depth map back-projection, and image feature matching. We develop a unified interface to integrate these methods and leverage the proposed CoorTrans module to align all data to the same FLU coordinate system. 3) Image Acquisition Interface: We integrate HTTP RESTful and TCP/IP protocols to form a unified image request interface, allowing image data to be obtained from any location with flexible resolutions and agent viewpoints.
3D Point Cloud Acquisition. For different simulators, we provide two methods to reconstruct the point cloud map of an entire scene. 1) Rasterized Sampling Reconstruction: For the UE5 + UnrealCV simulator, the MatrixCity [28] project offers a convenient rasterized sampling solution. We use the aforementioned lidar data acquisition interface to obtain the local point cloud at the sampling points. Since these data are already aligned within the same coordinate system, the point cloud map of the entire scene can be constructed by simply stitching local point clouds. For the UE4 + AirSim and GTA V simulators, we customize rasterized sampling points at appropriate resolutions, and perform sampling and reconstruction using the agent movement and lidar data acquisition interfaces. 2) Image-based Sparse Reconstruction: In 3D GS, the scene reconstruction process begins with the open-source COLMAP [43] framework, which geneoverlearates a sparse point cloud from input images. We directly export and use the point clouds obtained from this step.
Scene Semantic Segmentation. Vision-and-Language Navigation (VLN) requires meaningful landmarks as navigation targets. We perform semantic segmentation on four types of simulation scenes using the following three methods. 1) 3D Scene Understanding: A sequence of top-down views of the scene is captured in a rasterized format. We then use Octree graph [50] to extract 2D mask proposals, which are subsequently projected into the 3D point cloud space to generate semantic 3D segments. 2) Point Cloud Projection and Contour Extraction: We first acquire the point cloud of a scene, then project the voxelized point cloud onto a projection plane slightly above the ground. Using OpenCV, we perform a series of operations on the projected image, including binarization, erosion, and contour extraction, to obtain multiple instances along with their 2D coordinates. For each instance, the maximum height of the points within its neighborhood is used as the final height. This method provides a more computationally efficient yet coarser segmentation compared to the first approach, allowing users to choose based on their requirements. 3) Manual Annotation: When the point cloud quality of a scene is low or finer segmentation is required, we provide a method for annotating instances in the point cloud space by mouse clicks, based on ROS2 and RVIZ2. Users can annotate instances directly in the point cloud space using mouse clicks to define landmarks of interest for the task. This method is applicable to four simulators, i.e., UE + UnrealCV, UE + Airsim, GTA V, and 3D GS.
@@
Automatic Trajectory Generation. Leveraging the aforementioned point cloud map and segmentation tools, OpenFly can generate trajectories using the following two methods. 1) Path search based on customized action space: First, a global hash voxel map $M_{global}$ and a bird’s eye view (BEV) occupancy map $M_{bev}$ are constructed from the scene point cloud $P$. Second, the flight altitude is randomly selected within the user-defined height range, and landmarks that are not lower than the height threshold $H_{\tau}$ are chosen as targets. A starting point is selected within the distance range $[r,R]$ from the landmark, ensuring that it is not occupied in both $M_{global}$ and $M_{bev}$. Then, a point on the line connecting the starting point and the landmark, which is close to the landmark and unoccupied in $M_{bev}$, is chosen as the endpoint. Third, A collision-free trajectory from the starting point to the endpoint is generated using the A*  [16] pathfinding algorithm, where the granularity of exploration step size and direction can be adjusted according to the action space. Besides, by repeatedly selecting the endpoint as the new starting point, complex trajectories can be generated. Finally, utilizing OpenFly’s interface, images corresponding to the trajectory points can be obtained. 2) Path search based on grid: Google Map data does not allow image retrieval at arbitrary poses in the space. Thus we rasterize a pre-selected area and collect images from each grid point in all possible orientations. Starting and ending points are chosen within the grid points to generate trajectories. Corresponding images for these trajectory points are then selected from the pre-collected image set.
Automatic Instruction Generation. Previous research has predominantly relied on manual annotation to generate trajectory instruction, which is not only costly but also limits the scalability of datasets. To address this issue, we propose a highly automated language instruction generation method based on VLMs, e.g., GPT-4o. A straightforward method would be to submit all images to VLMs to analyze the trajectory and generate instructions. However, using all images introduces significant computational overhead and leads to information redundancy. For example, the ‘Forward’ action usually occupies a larger proportion of a flight trajectory, with ‘Turn Left/Turn Right’ or ‘Ascend/Descend’ actions taken when encountering key landmarks.
Based on the above findings, we split the complete trajectory into multiple sub-trajectories based on the occurrence of non-consistent actions, extracting key actions and images for processing and subsequent integration. Notably, slight angle adjustments often occur during flight to change the direction subtly, and a ‘slightly Turn Left/Right’ will be merged with subsequent ‘Forward’ actions. Specifically, suppose that the trajectory action sequence is [1, 1, 2, 1, 1, 2, 2, 1, 0], where 1, 2, and 0 denote ‘Forward’, ‘Turn Left’, and ‘Stop’, respectively. This trajectory would be split into four sub-trajectories, i.e., [1, 1], [2, 1, 1], [2, 2], and [1, 0]. The second sub-trajectory involves ‘slightly turn left’ and ‘move forward’. We submit the action sequence and the last captured image of each sub-trajectory to a VLM to generate descriptions of both action and landmarks. The sub-instructions are obtained similar to the following format:
  • ‘Move forward’ to ‘Landmark 1’.
  • ‘Slightly turn left and move forward’ to ‘Landmark 2’.
  • ‘Turn left’ towards ‘Landmark 3’.
  • ‘Move forward’ to ‘Landmark 4’.
These sub-instructions are then processed by a VLM/LLM again, where they are integrated into coherent and complete instructions. The detailed prompt used for the VLM, along with the complete responses, is provided in the supplementary material.
III-C Quality Control.
@@
At the same time, we built a data inspection platform to provide instructions, action sequences, and corresponding images to the examiners. If the instructions and trajectories align, they are considered qualified. We randomly select 3K samples from the entire dataset according to data distribution in Sec. V-A. After manually inspecting these samples, we find that the qualification rate reaches 91%.
IV Dataset Analysis
TABLE II: Comparison of different VLN datasets. Above the middle dividing line lies the ground-based datasets, while below is the aerial VLN datasets. $N_{traj}$: the number of total trajectories. $N_{vocab}$: vocabulary size. Path Len: the average length of trajectories, measured in meters. Intr Len: the average length of instructions. $N_{act}$: the average number of actions per trajectory.
Dataset $N_{traj}$ $N_{vocab}$ Path Len. Intr Len. Action Space $N_{act}$ Environment R2R [1] 7189 3.1K 10.0 29 graph-based 5 Matterport3D RxR [25] 13992 7.0K 14.9 129 graph-based 8 Matterport3D REVERIE [37] 7000 1.6K 10.0 18 graph-based 5 Matterport3D CVDN [45] 7415 4.4K 25.0 34 graph-based 7 Matterport3D TouchDown [7] 9326 5.0K 313.9 90 graph-based 35 Google Street View VLN-CE [23] 4475 4.3K 11.1 19 2 DoF 56 Matterport3D LANI [33] 6000 2.3K 17.3 57 2 DoF 116 CHALET ANDH [12] 6269 3.3K 144.7 89 3 DoF 7 xView AerialVLN [29] 8446 4.5K 661.8 83 4 DoF 204 AirSim + UE CityNav [26] 32637 6.6K 545 26 4 DoF - SensatUrban OpenUAV [47] 12149 10.8K 255 104 6 DoF 264 AirSim + UE Ours 100K 15.6K 99.1 59 4 DoF 35 AirSim + UE, GTA5 + Script Hook V, Google Earth Studio, 3D GS + SIBR viewers
(a) Difficulty level distribution.
(b) Length-height distribution.
(c) Action distribution with 1 type of ‘Forward’.
(d) Action distribution with 3 types of ‘Forward’.
Figure 4: Statistical analysis of trajectories.
IV-A Overview
@@
Using our toolchain, we collect 100k trajectories from 18 scenes, along with corresponding image sequences and language instructions. During the data generation process, we set a minimum motion step size of 3 meters to produce more granular trajectories. The details of our and previous VLN datasets are listed in Table. II, from which we can see that our dataset features a significantly larger number of trajectories and a more extensive vocabulary, as well as greater environmental diversity. In contrast, our average trajectory length and instruction length are relatively short. This is intentional, as we believe” short- and medium-range instructions are actually more in line with the usage habits of human users. This might be more beneficial for the aerial VLN field.
IV-B Trajectory Analysis
In addition to a rich variety of scenes, we also strive for diversity in the difficulty level, length, and height of the trajectory data. Based on the number of actions in one trajectory, we classify trajectories with fewer than 30 actions as ‘Easy’, those with the number of actions ranging from 30 to 60 as ‘Moderate’, and those with more than 60 actions as ‘Hard’. Fig. 4(a) shows the corresponding difficulty level distribution. Besides, compared with ground-based VLN, the aerial VLN task has more motion dimensions. Therefore, we set different trajectory lengths and flight heights to obtain rich data. Fig. 4(b) exhibits the distribution of these data, with their lengths ranging from 0 to 300 meters, and the heights ranging from 0 to 150 meters.
In the aerial VLN tasks of large-scale outdoor scenes, the proportion of moving forward is naturally higher than that of making adjustments in direction and altitude, as shown in Fig. 4(c). However, this highly unbalanced action distribution might cause the VLN model to overfit to the dominant action. To alleviate this problem, we divide the ‘Forward’ action into three granularities, i.e., 3m, 6m, and 9m. In the ground-truth trajectories, three consecutive ‘Forward’ actions will be combined into one ‘9m Forward’ action. At the end of a straight-moving trajectory, if the remaining distance is less than 9m, it will be combined into a ‘6m Forward’ action, or remain as a ‘3m Forward’ action. Fig. 4(d) presents the action distribution after this action merging process.
(a) Train set distribution.
(b) Test set distribution.
@@ IV-B Trajectory Analysis
Figure 5: The distribution of the data volume in different scenes under the Train and Test sets.
V OpenFly-Agent
V-A Dataset Split
Similar to previous works, we divide the dataset into three splits, i.e., Train, Test Seen, Test Unseen. Detailed data distributions are shown in Fig. 5. For the Train split, 7 scenes under the UE rendering engine account for $75.7\%$ of the total 100K data, since UE provides the largest number of scenes, where different amounts of trajectories are sampled according to the scenario area. The 4 scenes created by 3D GS are also the main part of the data, accounting for nearly $20\%$ of the total amount. To ensure visual quality, we only collect data from a high-altitude perspective using Google Earth, which accounts for $4.46\%$. The detailed information of the Test Seen and Test Unseen splits are as follows:
  • Test Seen: 1800 trajectories uniformly sampled from 11 previously seen UE and 3D GS scenarios.
  • Test Unseen: 1200 trajectories uniformly generated from 3 unseen scenarios, i.e., UE-smallcity, 3D GS-sjtu02, and a Los Angeles-like city in GTA V.
V-B Problem Definition
In the aerial VLN task, a UAV is randomly positioned within a 3D environment with its initial pose defined as $P=[x,y,z,\phi,\theta,\psi]$. At each timestamp $t$, the UAV perceives the surrounding environment through an egocentric image as its observation. Guided by natural language instructions, the task involves predicting the next navigation action. Notably, the UAV can utilize either the current observations or the frames from all previous timestamps to make its prediction.
V-C Model Architecture
As shown in Fig. 6, we take OpenVLA [22] as the baseline and design an end-to-end model for aerial VLN. In contrast, our model takes a sequence of images to indicate the observation instead of one image in the original OpenVLA. Moreover, to mitigate visual redundancy between adjacent video frames while maintaining key information, two strategies are proposed, i.e., keyframe selection and visual token merging. First, a series of candidate keyframes are selected. Then, these keyframes are merged temporally before and after the vision encoder, resulting in a compact sequence of visual tokens. Finally, the action decoder discretizes the predicted tokens into uniformly distributed bins, which are subsequently mapped to the 6 action types specific to drones.
V-C1 Keyframe Selection
The length of contextual visual tokens is a major challenge for VLMs when processing videos. Many open-source VLMs use uniform frame sampling [5, 39, 48] to reduce calculation, but this strategy is not suitable for aerial VLN, since it may miss frames containing key landmarks. To address this issue, we adopt a heuristic method to identify keyframes by detecting the change point of the UAV’s movement. We notice that sudden changes in the UAV’s trajectory are often caused by the observation of landmarks, which can serve as cues to determine keyframes. Specifically, we use the movement of the drone over time to draw turning curves, and the frames near the peaks of the wave are selected as candidate keyframes. The resulting data is interpolated and smoothed, forming a wave-like curve that represents the UAV’s movement.
To further ensure the precision of training data, scene segmentation maps collected in Sec. III-A are used on selected frames to detect key landmarks. Frames containing landmarks are selected as keyframes, yielding reasonably accurate results. Note that each sudden change of actions, e.g., from ‘Forward’ to ‘Turn Left’, will produce a set of keyframes. Consequently, we obtain several sets of keyframes for a long trajectory.
V-C2 Visual Token Merging
TABLE III: Comparison results on the test-seen split.

### v5

@@ Appendix A More Details of Rendering Engines and Data Resources
Table 3: Different 3D GS Scenes
@@
Appendix B Details of 3D GS Data Collection
From the UAV’s perspective, choosing the appropriate shooting altitude poses a dilemma, i.e., if the altitude is too low, the sparse point cloud generated during the initialization of the 3D GS reconstruction will be suboptimal, due to insufficient feature point matches between photos. In contrast, if the altitude is too high, the Gaussian reconstruction will result in an overly coarse training of details. After multiple attempts, the data collection plan using the M30T was determined as follows. For large-scale block scenes, oblique photography is performed at approximately twice the average building height using the default parameters of the M30T’s wide-angle camera, with a tilt angle of -45°. For landmark buildings with heights significantly different from the average height, additional targeted data collection is conducted at twice their height. This altitude setting can, to a certain extent, ensure both higher-quality point cloud initialization and Gaussian splatting training.
Appendix C Unified Interfaces
For data collection, we integrate all rendering engines and design three unified interfaces, i.e., the agent movement interface, the lidar data acquisition interface, and the image acquisition interface, allowing an agent to move and perceive the environment within any scene.
@@ Appendix C Unified Interfaces
  • Agent Movement Interface: We design a CoorTrans module, which implements a customized pose transformation matrix and scaling function to unify all coordinate systems into a meter-based FLU (Front-Left-Up) convention. This interface enables precise agent positioning, ensuring consistency and facilitating automatic trajectory generation.
  • Lidar Data Acquisition Interface: Lidar data is crucial for scene occupancy perception and essential for trajectory generation. Our platform supports different lidar data acquisition methods, including lidar sensor collection, depth map back-projection, and image feature matching. We develop a unified interface to integrate these methods and leverage the proposed CoorTrans module to align all data to the same FLU coordinate system.
  • Image Acquisition Interface: We integrate HTTP RESTful and TCP/IP protocols to form a unified image request interface, allowing image data to be obtained from any location with flexible resolutions and viewpoints.
Figure 6: Results of our point cloud acquisition and semantic segmentation. (a) Rasterized sampling point cloud reconstruction. (b) Image-based point cloud reconstruction. (c) Semantic 3D scene segmentation. (d) Point cloud projection and contour extraction.
Appendix D Results of Point Cloud Acquisition and Semantic Segmentation
This section presents results regarding the point cloud acquisition and scene semantic segmentation, as shown in Fig. 6.
Point Cloud Acquisition. OpenFly integrates data resources from different rendering engines. In order to obtain point cloud information from different scenes, we provide two point cloud acquisition methods. 1) For UE and GTAV scenes, we provide a tool that utilizes rasterized point cloud sampling and reconstruction to obtain a global point cloud. The results can be seen in Fig. 6(a). 2) For 3D GS scenes, we use COLMAP [36] to obtain relatively sparse point data from images, as shown in Fig. 6(b). Although the point cloud generated by this method is relatively sparse, it provides sufficient coverage information.
Scene Semantic Segmentation. To meet the requirements of different data resources and different segmentation granularities, our OpenFly offers three methods for obtaining 3D semantic segmentation of scenes. Here, we present results of segmentation methods other than manual annotation. 1) Fig. 6(c) illustrates the semantic segmentation results based on the off-the-shelf 3D scene understanding method Octree-Graph [37]. This method provides more granular results. 2) The result of semantic segmentation via point cloud projection and contour extraction is shown in Fig. 6(d). This method leverages high-precision point clouds to achieve instance segmentation for structures like buildings and trees, which directly contact the ground.
Appendix E A Comprehensive Introduction to Automatic Trajectory Generation
@@
Leveraging the point cloud map and segmentation tools, OpenFly offers two methods for trajectory generation for different scenes. 1) Path search based on customized action space: First, a global voxel map $M_{global}$ and a bird’s eye view (BEV) occupancy map $M_{bev}$ are constructed from the scene point cloud. Second, the flight altitude is randomly selected within the user-defined height range, and landmarks that are not lower than the height threshold $H_{\tau}$ are chosen as targets. A starting point is selected within the distance range $[r,R]$ from the landmark, ensuring that it is not occupied in both $M_{global}$ and $M_{bev}$. Then, a point on the line connecting the starting point and the landmark, which is close to the landmark and unoccupied in $M_{bev}$, is chosen as the endpoint. Third, A collision-free trajectory from the starting point to the endpoint is generated using the A*  [38] pathfinding algorithm, where the granularity of exploration step size and direction can be adjusted according to the action space. Besides, by repeatedly selecting the endpoint as the new starting point, complex trajectories can be generated. Finally, utilizing OpenFly’s interface, images corresponding to the trajectory points can be obtained. 2) Path search based on grid: Google Map data does not allow image retrieval at arbitrary poses in the space. Thus, we rasterize a pre-selected area and collect images from each grid point in all possible orientations. Starting and ending points are chosen within the grid points to generate trajectories. Corresponding images for these trajectory points are then selected from the pre-collected image set.
Appendix F Details of Instruction Generation
Except for the instruction generation described in the main paper. The remaining process is mainly divided into two parts: landmark feature extraction and sub-instruction fusion. A simplified prompt to the VLM and the corresponding response are probably like this.
  • Get Landmark features.
    System Prompt: You are an assistant who is proficient in image recognition. You can accurately identify the object in the picture and its characteristics that are different from the surrounding objects. I will give you the three final images you will see. Please focus on the last image and tell me the features of the target building and reply to me in the form of JSON.
    User: The target is the nearest prominent landmark to me. Answer me a dictionary like color:–, feature: –, size: –, type: –.
    GPT 4o: color: blue, feature: Steel, glass, size: medium size, type: building.
  • Instruction Fusion.
    System Prompt: You are an assistant proficient in text processing. You need to help me combine these scattered actions and landmarks into a sentence using words with similar meanings and more appropriate words, making them smooth, fluent, and accurate. If the landmarks of adjacent actions are similar or even identical, please use pronouns to refer to them.
    User: Multiple sub-instructions.
    GPT 4o: Move forward to a high-rise building with a noticeable logo at the top. Then, slightly turn left and go straight to a futuristic tower with a large spherical structure in the middle.
Appendix G Data Quality Control
@@
Manually Check. At the same time, we built a data inspection platform to provide instructions, action sequences, and corresponding images to human examiners. If an instruction describes all the actions and landmarks in a trajectory well, it is considered qualified. We randomly select 3K samples from the entire dataset according to the data distribution. After manually inspecting these samples, we find that they reach a high qualification rate of 91%. There is some ambiguity in the description of some landmarks in the remaining data, making it likely that these landmarks are not easily distinguishable from the surrounding environment. However, the examiners consider this not entirely unacceptable. In summary, most of the generated data feature good quality for the aerial VLN task.
Appendix H More Dataset Analyses
@@
Table 4 shows several commonly used statistical information of different VLN datasets, from which we can see that our dataset features a significantly larger number of trajectories and a more extensive vocabulary, as well as greater environmental diversity. In contrast, our average trajectory length and instruction length are relatively short. This is intentional, as we believe short- and medium-range instructions are actually more in line with the usage habits of human users. This might be more beneficial for the aerial VLN field.
Table 4: Comparison of different VLN datasets. $N_{traj}$: the number of total trajectories. $N_{vocab}$: vocabulary size. Path Len: the average length of trajectories, measured in meters. Intr Len: the average length of instructions. $N_{act}$: the average number of actions per trajectory.
Dataset $N_{traj}$ $N_{vocab}$ Path Len. Intr Len. Action Space $N_{act}$ Environment R2R [3] 7189 3.1K 10.0 29 graph-based 5 Matterport3D RxR [4] 13992 7.0K 14.9 129 graph-based 8 Matterport3D REVERIE [2] 7000 1.6K 10.0 18 graph-based 5 Matterport3D CVDN [5] 7415 4.4K 25.0 34 graph-based 7 Matterport3D TouchDown [1] 9326 5.0K 313.9 90 graph-based 35 Google Street View VLN-CE [6] 4475 4.3K 11.1 19 2 DoF 56 Matterport3D LANI [7] 6000 2.3K 17.3 57 2 DoF 116 CHALET ANDH [20] 6269 3.3K 144.7 89 3 DoF 7 xView AerialVLN [17] 8446 4.5K 661.8 83 4 DoF 204 AirSim + UE CityNav [21] 32637 6.6K 545 26 4 DoF - SensatUrban OpenUAV [18] 12149 10.8K 255 104 6 DoF 264 AirSim + UE Ours 100K 15.6K 99.1 59 4 DoF 35 AirSim + UE, GTA V, Google Earth Studio, 3D GS + SIBR viewers
@@
Figure 7: The distribution of the data volume in different scenes under the Train and Test sets. (a) Train set distribution. (b) Test set distribution.
Following previous studies [17, 20], we conducted a statistical analysis of the linguistic phenomena using 25 randomly selected instructions and compared the results with other VLN datasets, as detailed in Table 5. The analysis shows that the generated instructions exhibit rich linguistic phenomena such as ‘Reference’ and ‘Comparison’. Notably, our dataset is not the most complex one, since we believe that instructions in VLN tasks should be more aligned with real-life scenarios, rather than emphasizing length and complexity. This cognition is consistent with that of REVERIE [2]. Our instructions avoid overly lengthy and unrealistic expressions to some extent, making them more practical to command UAVs.
Fig. 7(a) shows the data distribution of the train set, where 7 UE scenes account for 75.7% of the total 100K data, 4 3D GS scenes account for nearly 20% of the total amount, and Google Earth data accounts for 4.46%. Fig. 7(b) presents the data distribution of the test set, where the seen data and unseen data account for 60% and 40%, respectively.
Table 5: Linguistic phenomena analysis using randomly selected 25 instructions. $p$ denotes the proportion of instructions showing the phenomenon, and $\mu$ represents the average number of times the phenomenon occurs in each instruction.
R2R[3] ANDH[20] AerialVLN[17] OpenFly
Phenomenon $p$ $\mu$ $p$ $\mu$ $p$ $\mu$ $p$ $\mu$ Example in OpenFly
Reference 100 3.7 92 1.9 100 9.7 100 2.7 …Advance to the large beige and brown building with windows
Coreference 32 0.5 8 0.1 68 1.8 52 1.4 …Continue moving forward to reach it
Comparison 4 0.0 32 0.4 20 0.2 60 0.7 …Move ahead to the medium-sized beige building…
Sequencing 16 0.2 8 0.1 68 3.7 64 0.8 …move forward to next large light brown building …
Allocentric Relation 20 0.2 32 0.4 56 4.6 76 1.0 …black large building featuring a billboard on its rooftop…
Egocentric Relation 80 1.2 32 0.4 100 7.1 100 2.4 … then slightly turn left as you move ahead towards…
Imperative 100 4.0 100 1.1 100 6.9 100 3.6 …. Proceed slightly straight and turn left ….
Direction 100 2.8 100 1.4 100 4.6 100 3.8 turn right and go ahead to …
Temporal Condition 28 0.4 20 0.2 76 5.6 72 0.9 …Continue straight until you reach …
Appendix I Implementation and Training Details